# Improving Expressivity of Graph Neural Networks using Localization

### v3

@@ Abstract
In this paper, we propose localized versions of Weisfeiler-Leman (WL) algorithms in an effort to both increase the expressivity, as well as decrease the computational overhead. We focus on the specific problem of subgraph counting and give localized versions of $k-$WL for any $k$. We analyze the power of Local $k-$WL and prove that it is more expressive than $k-$WL and at most as expressive as $(k+1)-$WL. We give a characterization of patterns whose count as a subgraph and induced subgraph are invariant if two graphs are Local $k-$WL equivalent. We also introduce two variants of $k-$WL: Layer $k-$WL and recursive $k-$WL. These methods are more time and space efficient than applying $k-$WL on the whole graph. We also propose a fragmentation technique that guarantees the exact count of all induced subgraphs of size at most 4 using just $1-$WL. The same idea can be extended further for larger patterns using $k>1$. We also compare the expressive power of Local $k-$WL with other GNN hierarchies and show that given a bound on the time complexity, our methods are more expressive than the ones mentioned in Papp and Wattenhofer (2022a).
11footnotetext: Indian Institute of Technology, Gandhinagar, India**footnotetext: Equal Contribution
@@ 1 Introduction
Graphs have been used for representing relational and structural data that appear in a variety of domains, ranging from social network analysis and combinatorial optimization to particle physics and protein folding Dill et al. (2008). In order to learn representations of these data for various downstream learning tasks such as graph classification, graph neural networks (GNNs) have emerged as very effective models. Given the various types of GNN-based models developed in recent years Kipf and Welling (2017); Veličković et al. (2018); Hamilton et al. (2018); Xu et al. (2019), researchers have attempted to characterize the expressive power of these models. Morris et al. (2019) showed the equivalence between message-passing GNNs and 1-Weisfeiler-Leman (WL) algorithm Weisfeiler and Leman (1968), which is a well known combinatorial technique for checking graph isomorphism and similarity. They also showed the equivalence between $k$-GNNs and $k-$WL. Thus, in this paper, by $k-$WL, we will be referring to the equivalent $k$-GNN model. In general, the expressiveness of $k-$WL, for any $k,$ is measured by its ability to identify non-isomorphic graphs and subgraphs. In this paper, we are using Folklore WL and state the results accordingly.
It has been shown that $(k+1)-$WL is more expressive than $k-$WL. The time and space complexity increases exponentially with $k$. Thus, it is infeasible to run $k-$WL on large graphs. Also, the $k-$WL hierarchy is crude as $3-$WL identifies almost all non-isomorphic graphs. Arvind et al. (2017) characterized the graphs that can be identified by $1-$WL. So, we are interested in coming up with a GNN hierarchy that can be easily extended without much computational overhead. More specifically, we want to define a GNN hierarchy whose expressiveness lies between $k-$WL and $(k+1)-$WL.
A count of specific patterns is very useful in determining the similarity between two graphs. However, detecting and counting the number of subgraphs is generally NP-complete as it is a generalization of the clique problem. There have been various works on efficient algorithms for some fixed patterns and restricted host graph classes Bressan (2018); Shervashidze et al. (2009); Bouritsas et al. (2020); Zhao et al. (2022); Shervashidze et al. (2011); Komarath et al. (2023); Ying et al. (2019); Liu et al. (2019). In Arvind et al. (2020) characterizes patterns whose count is invariant for $1-$WL and $2-$WL equivalent graphs. Also, there exists a GNN hierarchy, $S_{k}$ Papp and Wattenhofer (2022a), where each node has an attribute that counts the number of induced subgraphs of size at most $k$, that the node is participating in. It would be interesting to see whether a scalable GNN hierarchy exists, that is comparable to the $k-$WL and $S_{k}$ hierarchy. There also exists a GNN hierarchy, $M_{k},$ Papp and Wattenhofer (2022a); Huang et al. (2023)in which $k$ vertices are marked or deleted and a GNN model is run on the modified graph.
Various subgraph based GNN models have been proposed that tackle these questions Zhao et al. (2022); Zhang and Li (2021); Morris et al. (2018); Alvarez-Gonzalez et al. (2022); Maron et al. (2019); You et al. (2021); Frasca et al. (2022); Morris et al. (2021); Bevilacqua et al. (2021); Papp and Wattenhofer (2022b); Barceló et al. (2021); Huang et al. (2023). These GNNs have been effective in capturing more fine-grained patterns and relationships within the graphs, and are scalable for large graphs. Also, it has been shown that subgraph GNNs are more expressive than the traditional ones. Frasca et al. (2022) gave an upper bound on the expressive power of subgraph $1-$WL. This leads to the question of coming up with upper and lower bounds for the expressiveness of subgraph $k-$WL for arbitrary $k$.
Consider the task of counting the occurrence of $H$ as subgraph or induced subgraph in the host graph $G$. Given the effectiveness of subgraph $k-$WL methods in increasing the expressiveness of GNNs, we want to extend this method to the subgraph itself and check whether we can fragment the subgraph and learn the count of fragments of the subgraphs to get the actual count of $H$ in the graph. We are interested in evaluating its expressiveness in terms of subgraph and induced subgraph counting. Also, if there exists a GNN hierarchical model, we want to compare its expressiveness to pre-existing GNN hierarchies, as done in Papp and Wattenhofer (2022a).
@@ 1.1 Our Contributions
In this paper, we attempt to answer these questions. The main contributions of our work are listed below:
@@
    Characterizing expressiveness of Local $k-$WL: Given a graph $G=(V,E),$ we extract a $r$-hop subgraph rooted at each vertex $v\in V,$ say $G_{v}^{r},$ and run $k-$WL on $G_{v}^{r}.$ While GNNs based on subgraphs have been proposed in recent papers , this is the first work that gives both upper and lower bounds for the expressiveness of Local $k-$WL. We are also the first to characterize patterns that can be counted exactly by Local $k-$WL.
@@
    Layer $k-$WL: To improve the space and time complexity of Local $k-$WL, we propose the Layer $k-$WL method. For this method, instead of running $k-$WL on $G_{v}^{r}$, we run it on two consecutive layers of vertices. Here, the $i$th layer of vertices refers to the vertices that appear at an $i$-hop distance from $v$(or the $i$th layer breadth-first search(BFS)).
@@
    Recursive WL : Recursive WL is an alternative to $k-$WL. In this method, we first run $1-$WL to get a partitioning of vertices. Then we run $(k-1)-$WL on the vertices of each partition separately. It can be shown that this method is more expressive than $(k-1)-$WL and less expressive than $k-$WL. Also, since we are running $(k-1)-$WL on a smaller set of vertices, it has better space and time complexity than running $k-$WL.
@@
    Fragmentation : For the counting task, based on the pattern $H$ to be counted, the subgraph $G_{v}^{r}$ is further decomposed into simpler patterns for which the exact counts of subpatterns are already known. Thus, we need to learn the easier tasks in the subgraphs of $G_{v}^{r}$. So, a smaller $k$ is sufficient to count the number of patterns. Using this method, we show that all the patterns appearing as induced subgraphs of size four can be counted using just $1-$WL. This technique can be useful for counting larger and more complicated patterns. Thus, instead of training a GNN for the larger subgraph, we can train GNN models for the smaller patterns for counting and then combine their counts to get the count of the larger subgraph. Using the fragmentation technique, we use the model learned for predicting $K_{3}$ or a triangle to predict the number of $K_{4}$ in the graph. Similarly, if we have a model that can predict $K_{n}$ in a graph, then we can use it to predict $K_{n+1}$. In other words, we can reduce the counting of $K_{n+1}$ to a triangle counting problem with a minimal increase in the number of parameters.
@@
    Comparison with other GNN models : Papp and Wattenhofer (2022a) shows an analysis of four GNN models. We do a similar analysis for our models and compare them with the models mentioned in Papp and Wattenhofer (2022a). We show that our models are more expressive than the ones presented in that paper.
@@
In Section 2, we introduce some of the terms used throughout the paper. In Section 4, we introduce the localized variants of the $k-$WL algorithm and analyze their space and time complexities. In Section 5, we give theorems that characterize the expressiveness of the localized $k-$WL variants proposed in our work. In Section 6, we characterize the expressiveness of our methods in terms of subgraph and induced subgraph counting. We also discuss how to count the occurrences of $H$ in $G,$ using localized algorithms. We discuss the fragmentation approach in Section 7, followed by a theoretical comparison of GNN models in Section 8. The model architecture, along with the parameters used for the experiment, is explained in Section 9 We report the results of our experiments in Section 10 and conclude the paper with Section 11.
@@ 2 Preliminaries
We consider a simple graph $G(V,E)$. For basic definitions of graph theory, we refer the reader to West et al. (2001). The neighbourhood of a vertex $v$ is the set of all vertices adjacent to it in $G$ (denoted as $N_{G}(v)$). The closed neighbourhood of $v$ is the set of all neighbours, including the vertex $v$ (denoted as $N_{G}[v]$). A graph whose all the vertices are of same degree are called regular graph.
A graph $H$ is called a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. The subgraph induced on $S\subseteq V(G)$ is a graph whose vertex set $S$ contains all the edges in $G$ whose endpoints are in $S$ and is denoted by $G[S]$. The induced subgraph on a $r$-hop neighbourhood around vertex $v$ is denoted by $G_{v}^{r}$. Attributed subgraphs are coloured subgraphs, also referred to as Motifs.
The maximum distance from a vertex to all other vertices is called the eccentricity of the vertex. The minimum of the eccentricity over all the vertices is called the radius of the graph. The center of the graph is set of vertices such that eccentricity is minimum. For pattern counting, we pick one of the center vertex and call it key vertex.
@@
Homomorphism from graph $H$ to $G$ is a function from $V(H)$ to $V(G)$ such that if $\{u,v\}\in E(H)$ then $\{f(u),f(v)\}\in E(G)$. Given a pattern $H$, the set of all of its homomorphic images is called spasm of $H$. Two graphs $G$ and $H$ are isomorphic if there exists a bijective function $f:V(G)\xrightarrow{}V(H)$ such that $\{u,v\}\in E(G)$ if and only if $\{f(u),f(v)\}\in E(H)$. The orbit of a vertex $v$ in $G$ is the set of vertices to which $v$ can be mapped, and that mapping can be extended to automorphism (denoted by $Orbit_{G}(v)$).
We mention some of the structural parameters of graphs. Problems on graph with bounded structural parameters can be solved efficiently on bounded graph parameters.
Graph Parameters:
We first define tree decomposition of the graph as:
Given a graph $G$, we decompose it into tree structures, say $T,$ where the set of vertices of $T$ is a subset of set of vertices of $G.$ This decompostion has to satisfy the following constraints:
  1. 1.
    Every vertex of $G$ must lie in some bag associated with a vertex of $T$.
  2. 2.
    For each edge $\{v_{i},v_{j}\}$, there exists a bag containing having $v_{i}$ and $v_{j}$.
  3. 3.
    If a vertex $v_{i}\in V(G)$ belongs to two bags $B_{i}$ and $B_{j}$ associated with two vertices $u_{i}$ and $u_{j}$ of $T,$ then $v_{i}$ must be present in all the bags associated with the vertices belonging to the path connecting $u_{i}$ and $u_{j}$ in $T.$
The width of the tree decomposition is the maximum size of the bag minus one. The treewidth of the graph $G,$ $tw(G)$ is the minimum over all such decompositions. It is NP-hard to compute tree-width of graphs. However, there exists an efficient algorithm that checks for a fixed $k$, whether $tw(G)$ is at most $k$ Korhonen (2022); Korhonen and Lokshtanov (2022). Graph of bounded treewidth implies sparse graphs. However, for some sparse graph, the treewidth is unbounded. For example, grid graph on $n$ vertices has treewidth $\sqrt{n}$. The maximum of the treewidth over all its homomorphic images is called the hereditary treewidth of pattern $H,$ denoted by $htw(H)$.
Planar graphs are graphs that are $k_{5}$ and $k_{3,3}$ minor free. One can also say the graph that can be redrawn such that no edges cross each other are planar graphs. Also, the number of edges can be at most linear in the number of vertices. The Euler genus of a graph is defined in a similar manner. The genus of a graph is the minimum number such that the graph can be drawn on circle without intersecting edges.
Now, we look at graph classes that are dense but have nice structure such as complete graphs. Clique width has been defined for dense graphs. However, there does not exist efficient algorithm to check whether clique width of the given graphs is $k$ for $k\geq 4$. Rankwidth has been defined by Robertson and Seymour to handle dense graph classes. Given a graph $G$, to construct rankwidth decomposition, we define subcubic tree $T$. Given a bijection from the set of leaves of a tree to the set of vertices of $G$, we can construct rankwidth decomposition with respect to that bijection. The parent vertices in the tree contains union of vertices belonging to its two children. Note that deletion of single edge disconnects the tree and vertices of graph get partitioned into two subparts say $X$ and $V(G)\setminus X$. We define submatrix of adjacency matrix $A(X,V(G)\setminus X)$ where $a_{i,j}=1$ if and only if $i\in X$ and $j\in V(G)\setminus X$ are adjacent. Also, it has been shown that bounded clique width means bounded rank-width and vice versa Oum (2017).
@@ 3 Weisfeiler Leman Algorithm
Weisfeiler-Leman (WL) is a well known combinatorial algorithm that has many theoretical and practical applications. Color refinement($1-$WL) was first introduced in 1965 in Morgan (1965). The algorithm goes as follows:
@@
  • In the next iteration $i,$ we color the vertices by looking at the number of colors of vertices adjacent to each vertex $v,$ in the $(i-1)$th iteration, as
@@
  • The color class of any vertex $v\in V(G)$ can appear at most $1+\log n$ times and the running time is $\mathcal{O}(n^{2}\log n)$ Immerman and Sengupta (2019).
@@
The same idea was extended by Weisfeiler and Leman in which instead of coloring vertex, they colored all the two tuples based on edge, non-edge and $(v,v)$. In later iteration, the color gets refined for each two tuples based on their neighbourhood and common neighbourhood. This partition the set of two tuples of vertices. The iteration in which no further partition is being done are called stable coloring. Weisfeiler Leman algorithm which is known as $2-$WL algorithm.
@@
Let $\vec{x}=(x_{1},...,x_{k})\in V^{k}$,$y\in V$ , and $1\leq j\leq k$. Then, let $x[j,y]\in V^{k}$ denote the $k$-tuple obtained from $x$ by replacing $x_{j}$ by $y$. The $k$-tuples $\vec{x}[j,y]$ and $\vec{x}$ are said to be $j$-neighbors for any $y\in V$ We also say $\vec{x}[j,y]$ is the j-neighbor of $\vec{x}$ corresponding to $y$.
@@
  • The color class of any vertex $\vec{v}\in V^{k}$ can appear at most $\mathcal{O}(k\log n)$ times and running time is $\mathcal{O}(k^{2}n^{k+1}\log n)$ Immerman and Sengupta (2019).
@@
Color refinement ($1-$WL) can recognise almost all graphs Babai et al. (1980), while $2-$WL can recognise almost all regular graphs Bollobás (1982). The power of $WL$ increases with an increase in the value of $k$. The power of $k-$WL to distinguish two given graphs is same as with counting logic $C^{k+1}$ with $(k+1)$-variable. Also, the power of $k-$WL to distinguish two non-isomorphic graphs is equivalent to spoiler’s winning condition in $(k+1)$-bijective pebble game. Recently, Dell et al. (2018) has shown that the expressive power of $k-$WL is captured by homomorphism count. It has been shown that $G_{1}\simeq_{k}G_{2}$ if and only if $Hom(T,G_{1})=Hom(T,G_{2}),$ for all graphs $T$ of treewidth at most $k$.
The graphs that are identified by $1-$WL are Amenable graphs. There is a complete characterization of the amenable graphs in Arvind et al. (2017); Kiefer et al. (2015). In the original algorithm, we initially color all the vertices with color $1$. However, if we are given a colored graph as input, we start with the given colors as the initial colors. Also, we can color the edges, and run $1-$WL Kiefer et al. (2015).
Even if ${k-WL}$ may not distinguish two non-isomorphic graphs, two ${k-WL}$ equivalent graphs have many invariant properties. It is well known that two ${1-WL}$ equivalent graphs have the same maximum eigenvalue. Two graphs that are ${2-WL}$ equivalent are co-spectral and have the same diameter. Recently, V. Arvind et al. have shown the invariance in terms of subgraph existence and counts Arvind et al. (2020). They show the complete characterization of subgraphs whose count and existence are invariant for ${1-WL}$ equivalent graph pairs. They also listed down matching , cycles and path count invariance for ${2-WL}$. Also, there is a relation between homomorphism count and subgraph count Curticapean et al. (2017). The count of subgraphs is a function of the number of homomorphism from set of all homomorphic image of patterns. Hereditary treewidth of graph is defined as maximum of treewidth over all homomorphic images. So, if two graphs are ${k-WL}$ equivalent, then the count of all subgraphs whose $htw(G)$ is almost $k$ are same . However, running ${k-WL}$ takes $O(k^{2}\cdot n^{k+1}logn)$ time and $O(n^{k})$ space Immerman and Sengupta (2019). So, it is not practically feasible to run ${k-WL},$ for large $k,$ for graphs .
The expressive power of $k-$WL is equivalent to first order logic on $(k+1)$ variables with a counting quantifier. Let $G=(V,E),$ where $V$ is a set of vertices and $E$ is a set of edges. In logic, we define $V$ as the universe and $E$ as a binary relation. In Cai et al. (1992), they have proved that the power to distinguish two non-isomorphic graphs using $k-$WL is equivalent to $C^{k+1},$ where $C^{k+1}$ represents first order logic on $(k+1)$ variables with counting quantifiers (stated in Theorem 1). To prove this, they define a bijective $k$-pebble game, whose power is equivalent to $C^{k}$.
@@
The bijective k-Pebble game ($BP_{k}(G,H)$) has been discussed in Kiefer (2020); Cai et al. (1992); Grohe and Neuen (2021). Let graphs $G$ and $H$ have the same number of vertices and $k\in\mathbb{N}$. Let $v_{i},v\in V(G)$ and $w_{i},w\in V(H)$.
@@
The position of the game in bijective pebble game is the tuples of the vertices where the pebbles are placed.

### v5

# DiffCoALG: Differentiable Learning of Combinatorial Algorithms

@@ Abstract
Subgraph counting is a fundamental task for analyzing structural patterns in graph-structured data, with important applications in domains such as computational biology and social network analysis, where recurring motifs reveal functional and organizational properties. In this paper, we propose localized versions of the Weisfeiler-Leman (WL) algorithms to improve both expressivity and computational efficiency for this task. We introduce Local $k$-WL, which we prove to be more expressive than $k$-WL and at most as expressive as $(k+1)$-WL, and provide a characterization of patterns whose subgraph and induced subgraph counts are invariant under Local $k$-WL equivalence. To enhance scalability, we present two variants—Layer $k$-WL and Recursive $k$-WL—that achieve greater time and space efficiency compared to applying $k$-WL on the entire graph. Additionally, we propose a novel fragmentation technique that decomposes complex subgraphs into simpler subpatterns, enabling the exact count of all induced subgraphs of size at most $4$ using only $1$-WL, with extensions possible for larger patterns when $k>1$. Building on these ideas, we develop a three-stage differentiable learning framework that combines subpattern counts to compute counts of more complex motifs, bridging combinatorial algorithm design with machine learning approaches. We also compare the expressive power of Local $k$-WL with existing GNN hierarchies and demonstrate that, under bounded time complexity, our methods are more expressive than prior approaches.
@@ 1 Introduction
Graphs naturally model relational and structural data that arise across diverse domains, from social network analysis and combinatorial optimization to computational biology, particle physics, and protein folding (dill2008protein). Graph Neural Networks (GNNs) have emerged as powerful methods for learning graph representations for downstream tasks such as classification and link prediction. The expressive power of GNNs has been extensively studied, with message-passing GNNs shown to be equivalent to the $1$-dimensional Weisfeiler–Leman (WL) test (morris2019wl), a classical combinatorial tool for graph isomorphism. More generally, $k$-GNNs have been shown to correspond to the $k$-WL test (weisfeiler1968reduction), and thus, throughout this paper, we use $k$-WL interchangeably to denote the corresponding $k$-GNN hierarchy. In practice, the expressiveness of $k$-WL is determined by its ability to distinguish non-isomorphic graphs and capture subgraph patterns. However, while $(k+1)$-WL is provably more expressive than $k$-WL, both the time and space complexity of these algorithms grow exponentially in $k$, making higher-order WL infeasible in large-scale settings.
A central motivation driving this work lies in the task of subgraph counting, which is central to understanding complex graph structures. Counting specific subgraphs provides insights into graph similarity and functionality: in computational biology, this includes motifs critical for protein binding and molecular discovery; in social networks, structures such as triangles and stars reveal community dynamics and information propagation. Unfortunately, detecting and counting subgraphs — a generalization of the clique problem — is NP-complete in general. Prior works have therefore focused on specific families of patterns or restricted graph classes (shervashidze2009efficient; JMLR:v12:shervashidze11a; arvind2017graph; https://doi.org/10.48550/arxiv.1805.02089; https://doi.org/10.48550/arxiv.2006.09252; DBLP:conf/stacs/KomarathKMS23). At the same time, research has sought to enrich GNN hierarchies, including $S_{k}$-GNNs (papp2022theory), which augment nodes with counts of induced subgraphs of size up to $k$, and $M_{k}$-GNNs (papp2022theory; huang2023boosting), which modify graphs via marked or deleted vertices.
Traditional GNNs, being limited to the power of $1$-WL, cannot count even simple motifs such as triangles or cycles beyond length three (xu2019powerfulgraphneuralnetworks; morris2021weisfeilerlemanneuralhigherorder). Subgraph-based GNN models (zhao2022from; zhang2021nested; frasca) have emerged as extensions that apply GNNs on extracted local subgraphs or multiple partitions of the input graph, thereby improving expressiveness. While these approaches capture more fine-grained structures, they suffer from scalability concerns since GNNs must be run repeatedly over many subgraphs, creating prohibitive memory and time overhead.
To overcome these limitations, we propose localized variants of Weisfeiler–Leman algorithms that balance expressiveness and computational efficiency. Instead of running $k$-WL across the entire graph, we restrict its application to neighborhoods, resulting in the Local $k$-WL. Complementary to this, we introduce a novel fragmentation technique, which decomposes complex motifs into simpler subpatterns whose counts can be reliably obtained. Remarkably, we prove that all induced subgraphs of size at most 4 can be counted exactly using only $1$-WL, with natural extensions to larger motifs for higher $k$.
In addition, we design a differentiable three-stage learning framework that operationalizes this fragmentation principle: (1) identify required subpatterns, (2) count them using Local $k$-WL, and (3) aggregate counts into the global motif count. This approach provides both theoretical guarantees and compatibility with gradient-based learning, bridging combinatorial counting algorithms with machine learning paradigms. Finally, we analyze the expressive power of our methods relative to existing GNN hierarchies, demonstrating that under bounded complexity, our models achieve strictly higher expressivity.
@@ 1.1 Our Contributions
We summarize our key contributions:
@@
    Expressiveness of Local $k$-WL: We formally introduce Local $k$-WL, which applies $k$-WL to $r$-hop neighborhoods $G_{v}^{r}$ rooted at $v\in V.$ We provide, for the first time, both upper and lower bounds for its expressiveness and characterize precisely the subgraph patterns that can be distinguished and counted.
@@
    Layer $k$-WL: A scalable variant where $k$-WL is restricted to pairs of consecutive BFS layers, reducing time and space overhead compared to standard Local $k$-WL.
@@
    Recursive WL: A hierarchical alternative to $k$-WL. We first apply $1$-WL to partition nodes, and then run (k1)(k-1)-WL on smaller vertex sets. This achieves higher expressiveness than (k1)(k-1)-WL while maintaining better scalability than $k$-WL.
@@
    Fragmentation for Subgraph Counting: We propose a decomposition technique that reduces the counting of complex motifs to that of simpler subpatterns. For example, we reduce $K_{4}$ counting to repeated triangle counting. We prove that all induced subgraphs of size at most 4 can be counted with $1$-WL and extend this methodology to larger patterns.
@@
    Comparative Expressiveness: We provide a rigorous comparison of Local $k$-WL, Layer $k$-WL, and Recursive WL against existing GNN hierarchies such as $S_{k}$ and $M_{k}$. Our results show that, under comparable complexity bounds, our models are strictly more expressive.
In summary, our work advances subgraph counting in GNNs by combining the theoretical rigor of WL with scalable localized techniques, while introducing a new fragmentation paradigm that makes counting feasible, differentiable, and efficient for real-world graph applications.
@@
Section 2 introduces some of the terms used throughout the paper. In Section 4, we introduce the localized variants of the $k$-WL algorithm and analyze their space and time complexities. Section 5 gives theorems that characterize the expressiveness of the localized $k$-WL variants proposed in our work. In Section 6, we characterize the expressiveness of our methods in terms of subgraph and induced subgraph counting. We also discuss how to count the occurrences of $H$ in $G,$ using localized algorithms. We discuss the fragmentation approach in Section 7.4, followed by a theoretical comparison of GNN models in Section 8. The model architecture in Section 7. Discusson on implementation details, hyperparameters and results of our experiments is in Section 9 and conclude the paper with Section 10.
@@ 2 Preliminaries
We consider a simple undirected graph $G=(V,E)$ with vertex set $V$ and edge set $E$. For basic graph-theoretic definitions, we refer the reader to (west2001introduction). The neighbourhood of a vertex $v\in V$ is the set of all adjacent vertices, denoted by $N_{G}(v)$ (or $\mathcal{N}_{G}(v)$). The closed neighbourhood, denoted $N_{G}[v]$, includes $v$ itself. The degree of a node $v$ is denoted by $d_{v}$. The maximum distance from $v$ to any other vertex is its eccentricity, and the minimum eccentricity over all vertices gives the radius of the graph. A graph in which every vertex has the same degree is called a regular graph.
A graph $H$ is a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. The subgraph induced on $S\subseteq V(G)$, denoted $G[S]$, contains vertex set $S$ with all edges in $G$ whose endpoints lie in $S$. The induced subgraph on an $r$-hop neighbourhood of a vertex $v$ is denoted by $G_{v}^{r}$. The hop parameter $r$ depends on the pattern being counted. For example, $r=1$ suffices for counting triangles, while $r=2$ is required for counting $C_{4}$. Attributed subgraphs (or motifs) are subgraphs whose vertices or edges carry additional labels or colours.
@@
A graph homomorphism from $H$ to $G$ is a function $f:V(H)\to V(G)$ such that $\{u,v\}\in E(H)$ implies $\{f(u),f(v)\}\in E(G)$. The set of all homomorphic images of a pattern $H$ is called the spasm of $H$. Two graphs $G$ and $H$ are isomorphic if there exists a bijection $f:V(G)\to V(H)$ preserving adjacency. The orbit of a vertex $v$ in $G$, denoted $Orbit_{G}(v)$, is the set of vertices to which $v$ can be mapped under automorphisms of $G$.
2.1 Graph Parameters
Many hard graph problems become tractable on restricted graph classes characterized by structural parameters. We briefly review key notions.
Treewidth. A tree decomposition of $G$ expresses it as a tree of vertex “bags” satisfying: (i) every vertex of $G$ appears in some bag, (ii) every edge is contained in some bag, and (iii) for each vertex $v$, bags containing $v$ form a connected subtree. The width of a decomposition is the maximum bag size minus one, and the treewidth $tw(G)$ is the minimum width across all decompositions. Computing $tw(G)$ is NP-hard, but fixed-parameter algorithms exist (korhonen2022single; korhonen2022improved). The hereditary treewidth $htw(H)$ of a pattern $H$ is the maximum treewidth among its homomorphic images.
Planar graphs and genus. A graph is planar if it can be drawn without crossing edges, or equivalently, if it excludes $K_{5}$ and $K_{3,3}$ as minors. More generally, the Euler genus denotes the minimal surface on which a graph can be embedded without edge crossings.
Clique-width and rank-width. For dense graphs, clique-width provides a measure of complexity, but recognizing graphs of clique-width at most $k\geq 4$ is NP-hard. Rank-width, introduced by Robertson and Seymour, offers an alternative based on rank decompositions of the adjacency matrix. It is known that bounded clique-width implies bounded rank-width, and vice versa (oum2017rank).
2.2 Weisfeiler–Leman and Local $k$-WL
The Weisfeiler–Leman (WL) test (weisfeiler1968reduction) is a colour refinement algorithm for testing graph isomorphism. Its $k$-dimensional extension, $k$-WL, refines colours of $k$-tuples of vertices, with expressiveness increasing in $k$. Applying $k$-WL to a full graph with $n$ vertices requires $O(n^{k+1}\log n)$ time. In contrast, the Local $k$-WL algorithm applies $k$-WL to restricted neighbourhoods $G_{v}^{r}$, reducing complexity. For graphs of maximum degree $d$, local $k$-WL can be run in $O(n\cdot d^{r(k+1)}\log d)$ time.
2.3 Graph Neural Networks
Graph Neural Networks (GNNs) (DBLP:journals/corr/KipfW16) generalize neural architectures to graphs using recursive MESSAGE and AGGREGATE functions. At the $\ell$-th iteration, node embeddings are updated as:
$X_{v}^{(\ell)}=\texttt{UPDATE}^{(\ell)}\left(X_{v}^{(\ell-1)},\;\texttt{AGGREGATE}^{(\ell)}\Big\{\texttt{MESSAGE}^{(\ell)}\big(X_{v}^{(\ell-1)},\,X_{u}^{(\ell-1)}\big)\;\big|\;u\in N_{G}(v)\Big\}\right).$ (1)
It is known that standard message-passing GNNs are bounded in expressiveness by $1$-WL, motivating higher-order and subgraph-based extensions.
@@ 3 Weisfeiler Leman Algorithm
Weisfeiler-Leman (WL) is a well-known combinatorial algorithm that has many theoretical and practical applications. Color refinement(11-WL) was first introduced in 1965 in (morgan1965generation). The algorithm goes as follows:
@@
  • In the next iteration $i,$ we color the vertices by looking at the number of colors of vertices adjacent to each vertex $v,$ in the (i1)(i-1)th iteration, as
@@
  • The color class of any vertex $v\in V(G)$ can appear at most $1+\log n$ times and the running time is $\mathcal{O}(n^{2}\log n)$ (immerman2019kdimensional).
@@
The same idea was extended by Weisfeiler and Leman in which instead of coloring vertex, they colored all the two tuples based on edge, non-edge and $(v,v)$. In later iteration, the color gets refined for each two tuples based on their neighbourhood and common neighbourhood. This partition the set of two tuples of vertices. The iteration in which no further partition is being done are called stable coloring. Weisfeiler Leman algorithm which is known as 22-WL algorithm.
@@
Let $\vec{x}=(x_{1},\dots,x_{k})\in V^{k}$$y\in V$ , and $1\leq j\leq k$. Then, let $x[j,y]\in V^{k}$ denote the $k$-tuple obtained from $x$ by replacing $x_{j}$ by $y$. The $k$-tuples $\vec{x}[j,y]$ and $\vec{x}$ are said to be $j$-neighbors for any $y\in V$. We also say $\vec{x}[j,y]$ is the j-neighbor of $\vec{x}$ corresponding to $y$.
@@
  • The color class of any vertex $\vec{v}\in V^{k}$ can appear at most $\mathcal{O}(k\log n)$ times and running time is $\mathcal{O}(k^{2}n^{k+1}\log n)$ (immerman2019kdimensional).
@@
Color refinement (11-WL) can recognise almost all graphs (babai1980random), while $2$-WL can recognise almost all regular graphs (bollobas1982distinguishing). The power of $WL$ increases with an increase in the value of $k$. The power of $k$-WL to distinguish two given graphs is same as with counting logic $C^{k+1}$ with $(k+1)$-variable. Also, the power of $k$-WL to distinguish two non-isomorphic graphs is equivalent to spoiler’s winning condition in $(k+1)$-bijective pebble game. Recently, (Dell2018LovszMW) has shown that the expressive power of $k$-WL is captured by homomorphism count. It has been shown that $G_{1}\simeq_{k}G_{2}$ if and only if $Hom(T,G_{1})=Hom(T,G_{2}),$ for all graphs $T$ of treewidth at most $k$.
The graphs that are identified by 11-WL are Amenable graphs. There is a complete characterization of the amenable graphs in (arvind2017graph; kiefer2015graphs). In the original algorithm, we initially colour all the vertices with colour $1$. However, if we are given a coloured graph as input, we start with the given colours as the initial colours. Also, we can color the edges and run $1$-WL (kiefer2015graphs).
Even if kWL{k-WL} may not distinguish two non-isomorphic graphs, two kWL{k-WL} equivalent graphs have many invariant properties. It is well known that two 1WL{1-WL} equivalent graphs have the same maximum eigenvalue. Two graphs that are 2WL{2-WL} equivalent are co-spectral and have the same diameter. Recently, V. Arvind et al. have shown the invariance in terms of subgraph existence and counts (arvind2020weisfeiler). They show the complete characterization of subgraphs whose count and existence are invariant for 1WL{1-WL} equivalent graph pairs. They also listed down matching, cycles, and path count invariance for 2WL{2-WL}. Also, there is a relation between homomorphism count and subgraph count (curticapean2017homomorphisms). The count of subgraphs is a function of the number of homomorphisms from a set of all homomorphic images of patterns. Hereditary treewidth of a graph is defined as the maximum treewidth over all homomorphic images. So, if two graphs are kWL{k-WL} equivalent, then the count of all subgraphs whose $htw(G)$ is almost $k$ is the same. However, running kWL{k-WL} takes $O(k^{2}\cdot n^{k+1}logn)$ time and $O(n^{k})$ space (immerman2019kdimensional). So, it is not practically feasible to run kWL{k-WL} for large $k$, for graphs.
The expressive power of $k$-WL is equivalent to first-order logic on $(k+1)$ variables with a counting quantifier. Let $G=(V,E)$ where $V$ is a set of vertices and $E$ is a set of edges. In logic, we define $V$ as the universe and $E$ as a binary relation. In (cai1992optimal), they have proved that the power to distinguish two non-isomorphic graphs using $k$-WL is equivalent to $C^{k+1}$, where $C^{k+1}$ represents first-order logic on $(k+1)$ variables with counting quantifiers (stated in 1). To prove this, they define a bijective $k$-pebble game, whose power is equivalent to $C^{k}$.
@@
The bijective k-Pebble game ($BP_{k}(G,H)$) has been discussed in (kiefer2020power; cai1992optimal; rankwidth_WL). Let graphs $G$ and $H$ have the same number of vertices and $k\in\mathbb{N}$. Let $v_{i},v\in V(G)$ and $w_{i},w\in V(H)$.
@@
The position of the game in a bijective pebble game is the tuples of the vertices where the pebbles are placed.